OpenStudy (anonymous):

a man throws a ball off the top of a building and records the height of the ball at different times as shown in the table: Height of the ball time (s) Height (feet) 0 46 1 63 2 48 3 1 a. Find a quadratic model for the data b. Use the model to estimate the height of the ball at 2.5 seconds c. What is the ball's maximum height?

OpenStudy (anonymous):

I am totally lost...

OpenStudy (anonymous):

\[y=-a(x-h)^2+k\] (1,63) is a Vertex and also a maximum value \[y=-a(x-1)^{2}+63\] (0,46) x=0, y=46 \[46=-a(0-1)^{2}+63\] \[48=-a+63\] \[a=63-46=17\] therefore the equation: \[y=-17(x-1)^{2}+63\] at time 2.5 seconds you plug x=2.5 in the equation the maximum value (1,63)

OpenStudy (anonymous):

My teacher told me the answers for these but I'm lost as to how to get the anwers. She said the answers were: a. y= -16x^2 + 33x +46, where x is the number of seconds after release and y is the height in feet. b. 28.5 feet c. about 63 feet